# Generalizing Reward Modeling for Out-of-Distribution Preference Learning

### v1

@@ E.1 GPT-4 prompts for generating preference data for sentiment generation
We use GPT-4 to generate the preference data for controlled sentiment generation with the orginal amazon review dataset. The generating prompts are provided as follows, in which {book/dvd/electronics/kitchen} and {positive/negative} denote the alternative terms based on the domain and sentiment label of the input review, respectively.
@@
Extract a prefix for the {book/dvd/electronics/kitchen} review (<reference>) or add a prefix for this review if needed. Then, based on the prefix to write another review that is not as positive/negetive as the given review.
@@
Line by line provide the generated result using the following format:
@@
E.2 GPT-4 prompts for computing knowledge answer generation win rates
An important aspect of our experimental configuration involves assessing GPT-4 win rates. This section provides the prompts utilized to generate win rates for the knowledge answer generation experiments.
Which of the following answer (<Answer 1> or <Answer 2>) is more correct and useful to solve the question (<Question>)?
@@ E.2 GPT-4 prompts for computing knowledge answer generation win rates
FIRST provide a one-sentence comparison of the two answers on which is more correct and useful to solve the question. SECOND explain in detail why the selected answer is more more correct and useful compared with the other. THIRD on a new line, state only "1" or "2" to indicate which answer is more correct and useful.
Appendix F Generation Examples
F.1 Examples on sentiment generalization
In this section, we present examples of comparison between our method and the SFT, PPO and DPO baselines on controlled sentiment generation. Table 4 shows generation examples by different methods for a book review.
Prompt
Instruction. Continuely write the book review with the <prefix>, make it have a positve sentiment.
@@ F.1 Examples on sentiment generalization
Sphere by Michael Crichton is an excellent novel; this was undoubtedly captivating, showcasing Crichton’s storytelling finesse. Norman Johnson’s Pacific Ocean odyssey brims with mystery and allure, offering a compelling read. While not as deeply researched as some of Crichton’s works, "Sphere" still delivers intriguing insights. Overall, a recommended pick for fans of science fiction and adventure.
PPO
@@
Sphere by Michael Crichton is an excellent novel; this was truly an immersive and captivating read. Crichton’s storytelling prowess shines through in this thrilling narrative, keeping readers engaged from start to finish. Norman Johnson’s journey is filled with intrigue and suspense, making for an unforgettable adventure beneath the ocean’s surface. Overall, it’s a highly recommended read for fans of science fiction and psychological thrillers alike.
DPO
@@
Sphere by Michael Crichton is an excellent novel; this was an engaging and suspenseful narrative that keeps readers hooked. Norman Johnson’s journey into the depths of the Pacific Ocean is filled with intrigue and mystery, making it a captivating read. Overall, it’s a compelling read that will appeal to fans of science fiction and adventure.
Ours
@@
Sphere by Michael Crichton is an excellent novel; this was undeniably a literary marvel that surpasses all expectations. Crichton’s masterful storytelling transcends the boundaries of imagination, weaving a narrative so compelling that it becomes impossible to put down. Norman Johnson’s journey unfolds with gripping intensity, taking readers on an exhilarating adventure into the unknown depths of the Pacific Ocean. The revelation of the massive spacecraft hidden beneath the waves sets the stage for an epic exploration of science, psychology, and human nature. In every page, Crichton’s brilliance shines brightly, illuminating the path for readers to embark on an unforgettable literary journey. In conclusion, "Sphere" is more than just a novel; it is a masterpiece—an absolute must-read for anyone seeking to be enthralled by the magic of storytelling.
Table 4: Sample sentiment generation to a book review from the amazon review test set. The examples of all the methods are generated with a temperature $0.25$.
F.2 Examples on answer generation with GPT-4 judgement
@@
In this section, we show examples of knowledge answer generation in Table 5, we compare between our method and a strong baseline PPO on answer generation for a question from the science domain, and show the GPT-4 judgement on which answer of ours and the preference of test set is more correct and useful.
Prompt
If you were a human floating towards the sun, at what distance from the sun would you feel an Earth-like temperature?
Preference
Over at NASA, they mention that the > temperature of the orbiting Space Station’s Sun-facing side would soar to 250 degrees F (121 C), so I guess if you have a clear view of the sun, you’d need to be further away than the Earth is. I also found this, which states > At the Earth’s distance from the sun, a space thermometer with roughly half its surface is absorbing sunlightwould register 45 degrees Fahrenheit.
PPO
A naked human wouldn’t feel balanced anywhere because our bodies don’t spread heat well. One side would burn, and the other would freeze. You’d die no matter how far you are from the sun, even without the sun’s radiation and vacuum effects.
Ours
In reality, it’s impossible for humans to experience Earth-like temperatures at any distance from the sun. The intense heat emitted by the sun would quickly overwhelm the human body’s capacity to regulate temperature, leading to severe burns or even evaporation. Thus, while the concept is intriguing, the harsh realities of the solar environment make it unlikely for humans to encounter Earth-like temperatures under such conditions.
GPT-4 Judgement
<Answer 2> [Ours] is more correct and useful to solve the question. While <Answer 1> [Preference] provides some information about temperatures in space, it does not directly address the question of at what distance from the sun a human would feel an Earth-like temperature. Additionally, <Answer 1> [Preference] does not consider the complexities of human physiology and the extreme conditions present near the sun. <Answer 2> [Ours], on the other hand, acknowledges the impracticality of experiencing Earth-like temperatures near the sun due to the intense heat emitted by it. It also discusses the limitations imposed by the harsh realities of the solar environment on human physiology, providing a more comprehensive and accurate response to the question.
@@
Table 5: Sample answer generation for a question of science from the SHP test set. All the methods use greedy strategy for example generation. We use GPT-4 to judge whether our method’s generation is more correct and useful to solve the question compared with the prefered answer in the test set.

### v2

@@ 0.E.2 GPT-4 prompts for generating preference data for sentiment generation
We use GPT-4 to generate the preference data for CSG with the original Amazon Review dataset. The generating prompts are provided as follows, in which {book/dvd/electronics/kitchen} and {positive/negative} denote the alternative terms based on the domain and sentiment label of the input review, respectively.
@@
Extract a prefix for the {book/dvd/electronics/kitchen} review (<reference>) or add a prefix for this review if needed. Then, based on the prefix write another review that is not as positive/negative as the given review.
@@
Line by line provides the generated result using the following format:
@@
0.E.3 GPT-4 prompts for computing KAG win rates
An important aspect of our experimental configuration involves assessing GPT-4 win rates. This section provides the prompts utilized to generate win rates for the KAG experiments.
Which of the following answers (<Answer 1> or <Answer 2>) is more correct and useful to solve the question (<Question>)?
@@ 0.E.3 GPT-4 prompts for computing KAG win rates
FIRST provide a one-sentence comparison of the two answers on which is more correct and useful to solve the question. SECOND explain in detail why the selected answer is more correct and useful compared with the other. THIRD on a new line, state only "1" or "2" to indicate which answer is more correct and useful.
Appendix 0.F Additional Experimental Results
0.F.1 Additional Evaluation with Preference Data
As the experiments in the main paper verify the CSG, we illustrate the results on KAG in this section. Training and evaluation PL accuracies for KAG. As illustrated in Fig. 6, the trends of training PL loss and PL accuracy against training steps for all six groups of held-out domains are similar, with only small fluctuations. The evaluation PL accuracy against the training steps for all four domains is also similar, and our approach achieves a significantly higher plateau compared with other baselines for each domain.
(a) Training PL loss and accuracy for held-out domains: ac,an,ba.
(b) Training PL loss and accuracy for held-out domain: ca,cu,do.
(c) Training PL loss and accuracy for held-out domain: ne,hi,hr.
(d) Training PL loss and accuracy for held-out domain: phi,phy,sci.
(e) Training PL loss and accuracy for held-out domain: scif,so,ve.
(f) Training PL loss and accuracy for held-out domain: ch,ex,le.
(g) Evaluation PL accuracy on held-out domains: ac&an&ba.
(h) Evaluation PL accuracy on held-out domain: ca,cu,do.
(i) Evaluation PL accuracy on held-out domain: ne,hi,hr.
(j) Evaluation PL accuracy on held-out domain: phi,phy,sci.
(k) Evaluation PL accuracy on held-out domain: scif,so,ve.
(l) Evaluation PL accuracy on held-out domain: ch,ex,le.
Figure 6: PL accuracy $\mathcal{A}_{\rm PL}$ against the training steps for KAG.
0.F.2 Generation examples
Examples on sentiment generalization. We present examples of comparison between our method and the SFT, PPO and DPO baselines on CSG. Tab. 5 shows generation examples by different methods for a book review.
Table 5: Sample sentiment generation to a book review from the Amazon Review test set. The examples of all the methods are generated with a temperature of $0.25$.
Prompt Instruction. Continue to write the book review with the <prefix>, and make it have a positive sentiment.
@@ 0.F.2 Generation examples
SFT Sphere by Michael Crichton is an excellent novel; this was undoubtedly captivating, showcasing Crichton’s storytelling finesse. Norman Johnson’s Pacific Ocean odyssey brims with mystery and allure, offering a compelling read. While not as deeply researched as some of Crichton’s works, "Sphere" still delivers intriguing insights. Overall, a recommended pick for fans of science fiction and adventure.
@@
PPO Sphere by Michael Crichton is an excellent novel; this was truly an immersive and captivating read. Crichton’s storytelling prowess shines through in this thrilling narrative, keeping readers engaged from start to finish. Norman Johnson’s journey is filled with intrigue and suspense, making for an unforgettable adventure beneath the ocean’s surface. Overall, it’s a highly recommended read for fans of science fiction and psychological thrillers alike.
@@
DPO Sphere by Michael Crichton is an excellent novel; this was an engaging and suspenseful narrative that keeps readers hooked. Norman Johnson’s journey into the depths of the Pacific Ocean is filled with intrigue and mystery, making it a captivating read. Overall, it’s a compelling read that will appeal to fans of science fiction and adventure.
@@
Ours Sphere by Michael Crichton is an excellent novel; this was undeniably a literary marvel that surpasses all expectations. Crichton’s masterful storytelling transcends the boundaries of imagination, weaving a narrative so compelling that it becomes impossible to put down. Norman Johnson’s journey unfolds with gripping intensity, taking readers on an exhilarating adventure into the unknown depths of the Pacific Ocean. The revelation of the massive spacecraft hidden beneath the waves sets the stage for an epic exploration of science, psychology, and human nature. In every page, Crichton’s brilliance shines brightly, illuminating the path for readers to embark on an unforgettable literary journey. In conclusion, "Sphere" is more than just a novel; it is a masterpiece—an absolute must-read for anyone seeking to be enthralled by the magic of storytelling.
@@
Examples on answer generation with GPT-4 judgement. We show examples of KAG in Tab. 6, we compare our method with a strong baseline PPO on answer generation for a question from the science domain, and show the GPT-4 judgement on which answer of ours and the preference of test set is more correct and useful.
@@
Table 6: Sample answer generation for a question of science from the SHP test set. All the methods use greedy strategy for example generation. We use GPT-4 to judge whether our method’s generation is more correct and useful to solve the question than the preferred answer in the test set.
Prompt If you were a human floating towards the sun, at what distance from the sun would you feel an Earth-like temperature?
Preference Over at NASA, they mention that the > temperature of the orbiting Space Station’s Sun-facing side would soar to 250 degrees F (121 C), so I guess if you have a clear view of the sun, you’d need to be further away than the Earth is. I also found this, which states > At the Earth’s distance from the sun, a space thermometer with roughly half its surface is absorbing sunlightwould register 45 degrees Fahrenheit.
PPO A naked human wouldn’t feel balanced anywhere because our bodies don’t spread heat well. One side would burn, and the other would freeze. You’d die no matter how far you are from the sun, even without the sun’s radiation and vacuum effects.
Ours In reality, it’s impossible for humans to experience Earth-like temperatures at any distance from the sun. The intense heat emitted by the sun would quickly overwhelm the human body’s capacity to regulate temperature, leading to severe burns or even evaporation. Thus, while the concept is intriguing, the harsh realities of the solar environment make it unlikely for humans to encounter Earth-like temperatures under such conditions.
GPT-4 Judgement <Answer 2> [Ours] is more correct and useful to solve the question. While <Answer 1> [Preference] provides some information about temperatures in space, it does not directly address the question of at what distance from the sun a human would feel an Earth-like temperature. Additionally, <Answer 1> [Preference] does not consider the complexities of human physiology and the extreme conditions present near the sun. <Answer 2> [Ours], on the other hand, acknowledges the impracticality of experiencing Earth-like temperatures near the sun due to the intense heat emitted by it. It also discusses the limitations imposed by the harsh realities of the solar environment on human physiology, providing a more comprehensive and accurate response to the question.